### v1

# A Novel $J_{M}$-Based Method for Interpolation in Metric Spaces

Abstract
We present a new interpolation method for metric spaces, termed the $J_{M}^{\prime}$-interpolated metric space. Building on previous work where two interpolation methods were developed in analogy with Peetre’s classical $K$ and $J$ methods, this approach introduces a distinct construction based on approximating sequences in the intersection space $X_{0}\cap X_{1}$ equipped with the metric $J_{M}(1)$. The method defines a relative completion within the union space $X_{0}\cup X_{1}$, yielding a well-defined interpolated metric space with controllable relationships between the $J_{M}$ and $K_{M}$ metrics. While we do not compare this method to the previously developed approaches nor explore its correspondence with classical interpolation functors in normed spaces, the construction enriches the set of tools available for metric-space interpolation theory. Furthermore, as expected, these methods still preserve the Lipschitz property of an operator between two interpolated spaces. Potential applications include general Nonlinear Analysis and Wasserstein-type spaces where the sequence-based nature of the method may provide practical advantages.
Introduction
In a previous work published in 2024 [see [13]], we introduced two methods of interpolation for general metric spaces, inspired by the classical $K$ and $J$ methods in the normed setting of Peetre [see [10] and [3]]. These constructions allowed the extension of classical interpolation ideas to the broader context of metric spaces, defining analogues of the $K$- and $J$-interpolation spaces and establishing a structured approach to study interpolation without relying on linearity or normed assumptions. This type of interpolation acts on metric spaces and preserves the Lipschitz continuity of operators in the interpolated spaces, in a manner analogous to interpolation functors in classical Banach space theory. The methods provided a framework for constructing interpolated metric spaces that capture intermediate structures between two given metric spaces, opening avenues for applications in areas where metric spaces arise naturally, such as nonlinear analysis, optimal transport, and metric geometry.
In this article, we develop a new interpolation method, the $J_{M}^{\prime}$-interpolated metric space, which complements the previously introduced methods by offering a fundamentally different construction. Unlike the prior methods, which relied on direct adaptations of the $K$ and $J$ functionals, this new method is based on approximating sequences within the intersection space $X_{0}\cap X_{1}$, endowed with the metric $J_{M}(1)$, and its relative completion inside the union space $X_{0}\cup X_{1}$. This sequence-based approach introduces a unique perspective, which could prove advantageous in situations where the behavior of sequences in the intersection space is critical or where the structure of the union space provides natural constraints.
The primary motivation behind proposing a new interpolation method is that expanding the toolbox of available interpolation techniques enriches the study of metric-space interpolation, allowing researchers to choose the most suitable method depending on the specific problem or application. This methodological diversity is particularly relevant for applications to Nonlinear Analysis and spaces with nontrivial geometries, such as Wasserstein-type spaces [see [1] and [15] and [16]], where the sequence-oriented nature of this approach might yield novel insights or simplifications.
It is important to note that the goal of this article is the construction of the new interpolation method itself. We do not investigate its relationship with the previously developed methods, nor do we attempt to determine whether it generates interpolation functors in normed spaces that are equivalent to classical constructions by Peetre or Lions. Our focus is on establishing the theoretical foundation of this new approach and analyzing its intrinsic properties, such as its metric structure, continuity of inclusions, and basic interpolation inequalities. As a final remark, we suggest that future research could explore potential applications of this method, particularly in Wasserstein-type spaces, where the specific sequence-based construction might reveal previously inaccessible properties of interpolated structures.
1 Preliminaries
All results and definitions in this section are found in the work [see [13]]. Thus, we decided to omit the proofs of basic lemmas of the theory that fall outside the scope of this text. The interested reader can verify the proofs in the reference above.
1.1 Relative Completion of Metric Spaces
Definition 1.1.
Let $(A,d_{A})$ and $(B,d_{B})$ be two metric spaces such that $A\subseteq B$ and there exists $c>0$ such that
$d_{B}(x,y)\leq c\cdot d_{A}(x,y),\quad x,y\in A.$
We define the relative completion of $(A,d_{A})$ in $(B,d_{B})$, denoted by $A_{B}$, by
$A_{B}:=\{x\in B;\,\,\textbf{$\exists$ Cauchy sequence $(x_{n})$ in $(A,d_{A})$ and $d_{B}(x_{n},x)\to 0$}\}.$
A sequence $(x_{n})$ as above is called an approximating sequence of $x$ in $(A,d_{A})$.
There are some intrinsic properties of the relative completion of metric spaces that we list below:
  1. 1.
    This completion can be metrized with a metric $d_{A,B}$ that satisfies
    $d_{B}(x,y)\leq c\cdot d_{A,B}(x,y)\leq d_{A}(x,y),\quad x,y\in A$
    and
    $d_{B}(x,y)\leq c\cdot d_{A,B}(x,y),\quad x,y\in A_{B}.$
    Thus, $A\subseteq A_{B}\subseteq B$ with continuous inclusions.
  2. 2.
    The set $A$ is dense in $(A_{B},d_{A,B})$.
  3. 3.
    If $(B,d_{B})$ is complete, then $(A_{B},d_{A,B})$ is also complete.
  4. 4.
    The set $A$ is dense in $(A_{B},d_{A,B})$.
  5. 5.
    The metrics $d_{A,B}$ and $d_{A}$ coincide on $A$.
1.2 Pairs of Compatible Metric Spaces
Definition 1.2.
Let $X_{0}$ and $X_{1}$ be two sets such that $X_{0}\cap X_{1}\neq\emptyset$ and both are subsets of a set $X$. Furthermore, assume that $d_{0},d_{1}$ and $d_{X}$ are metrics on the sets $X_{0}$, $X_{1}$ and $X$, respectively, such that for convenient positive constants $C_{0}$ and $C_{1}$ we have
$d_{X}(x,y)\leq C_{i}\cdot d_{i}(x,y),\quad\forall\,x,y\in X_{i},\quad i\in\{0,1\}.$ (1)
We call this object a pair of compatible metric spaces and denote it by $\overrightarrow{X}$ or by $(X_{0},X_{1})_{X}$.
Let $(X_{0},X_{1})_{X}$ and $t>0$. For $(x,y)\in X_{0}^{2}\cup X_{1}^{2}$, define the function $h_{t}$ which will be given by
$h_{t}(x,y)=\left\{\begin{array}[c]{l}\min\{d_{0}(x,y),t\cdot d_{1}(x,y)\},\text{ if }x,y\in X_{0}\cap X_{1};\\ d_{0}(x,y),\text{ if}\,\,\{x,y\}\subseteq X_{0}\setminus X_{1};\\ t\cdot d_{1}(x,y),\text{ if}\,\,\{x,y\}\subseteq X_{1}\setminus X_{0}.\end{array}\right$.
Definition 1.3.
Given $x,y\in X_{0}\cup X_{1}$, with $X_{0}\cap X_{1}\neq\emptyset$, an admissible linking sequence from $x$ to $y$ in $X_{0}\cup X_{1}$ is any finite sequence of points $x_{0},x_{1},\ldots,x_{n}\in X_{0}\cup X_{1}$ that satisfies:
  • $x_{0}=x$ and $x_{n}=y$;
  • For each $k\in\{0,1,\ldots,n-1\}$, $(x_{k},x_{k+1})\in X_{0}^{2}\ cupX_{1}^{2}$ holds.
Let us denote an admissible linking sequence $(x,x_{1},x_{2},\ldots,x_{n-1},y)$ from $x$ to $y$ in $X_{0}\cup X_{1}$ by $(x_{n}^{x,y})_{adm}$. When all $n$ points of the sequence belong to the same set $Z$, we will say that it is a linking sequence from $x$ to $y$ in $Z$ and denote it only by $(x_{n}^{x,y})$. The set of all linking sequences from $x$ to $y$ in $Z$ is denoted by $x,y]_{Z}$.
Remark 1.1.
Note that every admissible linking sequence is a linking sequence in the union $X_{0}\cup X_{1}$, but not every linking sequence in $X_{0}\cup X_{1}$ is admissible.
1.3 The $K_{M}$ Functional and its properties
Definition 1.4 (The $K_{M}$-Metric Functional).
Let $(X_{0},X_{1})_{X}$ be a pair of compatible metric spaces and $t>0$. Given $x,y\in X_{0}\cup X_{1}$, we define the $K_{M}$-metric functional by
$K_{M}(t;x,y):=\inf\,\sum_{k=0}^{n-1}h_{t}(x_{k},x_{k+1}),$ (2)
where the infimum is taken over all admissible linking sequences from $x$ to $y$ in the set $X_{0}\cup X_{1}$.
From now on, unless otherwise stated and when there is no possibility of confusion due to notation, $(X_{0},d_{0})$ and $(X_{1},d_{1})$ will be considered a pair of compatible metric spaces with respect to the space $(X,d_{X})$, that is, they will represent $(X_{0},X_{1})_{X}$.
Lemma 1.1.
[…]
Lemma 1.2.
If $a,b\in(0,\infty)$ and $x,y\in X_{0}\cup X_{1}$, then
$K_{M}(a;x,y)\leq\max\biggr\{1,\frac{a}{b}\biggr\}\,K_{M}(b;x,y)$ (3)
1.4 The $J_{M}$ Functional
Lemma 1.3.
If $x,y\in X_{0}\cap X_{1}$ and $t>0$, defining
$J_{M}(t;x,y):=\max\{d_{0}(x,y),t\cdot d_{1}(x,y)\},$
[…]
Lemma 1.4.
If $a,b\in(0,\infty)$ and $x,y\in X_{0}\cap X_{1}$, then
$J_{M}(a;x,y)\leq\max\biggr\{1,\frac{a}{b}\biggr\}\cdot J_{M}(b;x,y),\quad x,y\in X_{0}\cap X_{1};$ (4)
Lemma 1.5.
If $x,y\in X_{0}\cap X_{1}$ and $a,b\in(0,\infty)$, then
$K_{M}(a;x,y)\leq\min\biggr\{1,\frac{a}{b}\biggr\}\cdot J_{M}(b;x,y).$ (5)
2 The $K_{M}$-Metric Interpolation Space
In this section we will define and highlight the main properties of the interpolated metric space.
Definition 2.1.
Let $\theta\in(0,1)$ and $q\in[1,\infty]$. Given a sequence $(x_{k})_{k\in\ \mathbb{Z}}$ of real numbers, we define the functional $\Gamma_{\theta,q}$ by
$\Gamma_{\theta,q}((x_{k})_{k\in\mathbb{Z}}):=\begin{cases}\biggr[\sum_{k\in\mathbb{Z}}[2^{-k\theta}|x_{k}|]^{q}\biggr]^{1/q},&q<\infty\\ \sup_{k\in\mathbb{Z}}2^{-k\theta}|x_{k}|,&q=\infty.\end{cases}$ (6)
Lemma 2.1.
The functional
$\beta_{\theta,q}(x,y):=\Gamma_{\theta,q}((K_{M}(2^{k};x,y)_{k\in\mathbb{Z}})$ (7)
is a metric on the set $X_{0}\cap X_{1}$.
We will refer, for simplicity, to the metrics $J_{M}(t;\cdot,\cdot)$ and $K_{M}(t;\cdot,\cdot)$ as $J_{M}(t)$ and $K_{M}(t)$, respectively.
Definition 2.2 (The $K_{M}$-interpolated metric space).
The interpolated metric space by the classical $K_{M}$ method is denoted by $\overrightarrow{X}_{\theta,q}^{K_{M}}$ and is the relative completion of the metric space
$(X_{0}\cap X_{1},\beta_{\theta,q})$
in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. The metric of this space is denoted by $\mathcal{D}_{\theta,q}$ and equals $\mathcal{D}_{\theta,q}(x,y)=\beta_{\theta,q}(x,y)$, for all $x,y\in X_{0}\cap X_{1}$.
Lemma 2.2.
Given $\theta\in(0,1)$ and $q\geq 1$, we have
$\beta_{\theta,q}(x,y)\leq M_{\theta,q}\cdot J_{M}(1;x,y),\quad x,y\in X_{0}\cap X_{1}.$ (8)
And furthermore,
$M_{\theta,q}\cdot d_{X}(x,y)\leq\mathcal{D}_{\theta,q}(x,y),\quad x,y\in\overrightarrow{X}_{\theta,q}^{K_{M}}.$ (9)
where
$M_{\theta,q}=\displaystyle\Gamma_{\theta,q}(\min\{1,2^{k}\}_{k\in\mathbb{Z}})>0.$
3 The $J_{M}^{\prime}$-Interpolated Metric Space
Definition 3.1 (Bi-infinite Sequences).
A sequence is said to be bi-infinite when it is indexed in the set of integers. A bi-infinite sequence $(z_{k})_{k\in\mathbb{Z}}$ is said to converge to the point $z$ in a metric space $(M,d)$ if
$\lim_{n\to\infty}d(z_{n},z)=0\quad\text{and}\quad\lim_{n\to-\infty}d(z_{n},z)=0.$
We know that $X_{0}\cap X_{1}\subseteq X_{0}\cup X_{1}$ and that $K_{M}(1)$ is a metric on $X_{0}\cup X_{1}$. We will metrize the set $X_{0}\cap X_{1}$ with a metric satisfying the relative completion condition and then complete this space equipped with this metric, with respect to the space $(X_{0}\cup X_{1},K_{M}(1))$.
Given $x,y\in X_{0}\cap X_{1}$ and $(x_{n})$ and $(y_{n})$ approximating sequences of $x$ and $y$ in $(X_{0}\cap X_{1},J_{M}(1))$, respectively, let us define a new bi-infinite sequence, denoted by $(S_{k}^{x,y})_{k\in\mathbb{Z}}$, which depends on these approximating sequences and whose general term is given by:
$S_{k}^{x,y}:=\begin{cases}x_{k},&\text{if $k\geq 0$}\\ y_{-k},&\text{if $k<0$}\end{cases}$
Such a sequence will be called a bi-infinite linking sequence from $x$ to $y$ in $(X_{0}\cap X_{1},J_{M}(1))$. On $X_{0}\cap X_{1}$ let us define the functional $p_{\theta,q}$ given by
$p_{\theta,q}(x,y):=\inf\Gamma_{\theta,q}((J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y})),$ (10)
where the infimum is taken over all approximating sequences of $x$ and $y$.
Definition 3.2 (Separator Functional).
Let $M\neq\emptyset$ and $p:M\times M\to[0,\infty)$. We say that $p$ is a separator functional of points in $M$ and
$p(x,y)=0\Longleftrightarrow x=y.$
Theorem 3.1.
The functional $p_{\theta,q}$ is a separator on $X_{0}\cap X_{1}$.
Proof.
Let $x,y\in X_{0}\cap X_{1}$ and $(x_{n})$ and $(y_{n})$ be two arbitrary approximating sequences of $x$ and $y$ in $(X_{0}\cap X_{1},J_{M}(1))$ and let $(S_{j}^{x,y})_{j\in\mathbb{Z}}$ be the bi-infinite linking sequence from $x$ to $y$ in $(X_{0}\cap X_{1},J_{M}(1))$. Fixing $N\in\mathbb{N}$, we have
$K_{M}(1;S_{N}^{x,y},S_{-N}^{x,y})\leq\sum_{j=-N}^{N-1}K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y}).$
Thus, taking the $\lim\sup_{N}$, we have
$K_{M}(1;x,y)\leq\sum_{j\in\mathbb{Z}}K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y}).$
Now, using Lemma 1.5, with $a=1$ and $b=2^{j}$ in each term $K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y})$, we have
$K_{M}(1;x,y)\leq\sum_{j\in\mathbb{Z}}K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y})\leq\sum_{j\in\mathbb{Z}}\min\{1,1/2^{j}\}\cdot J_{M}(2^{j};S_{j}^{x,y},S_{j+1}^{x,y}).$
We can also write, for all $\theta\in(0,1)$:
$K_{M}(1;x,y)\leq\sum_{j\in\mathbb{Z}}[2^{j\theta}\cdot\min\{1,1/2^{j}\}]\cdot[2^{-j\theta}\cdot J_{M}(2^{j};S_{j}^{x,y},S_{j+1}^{x,y})].$
Using the Hölder Inequality in the last sum, we have
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot\Gamma_{\theta,q}((J_{M}(2^{j};S_{j}^{x,y},S_{j+1}^{x,y})_{j\in\mathbb{Z}}),$ (11)
where
$M_{\theta,q}:=\biggr[\sum_{j\in\mathbb{Z}}[2^{j\theta}\cdot\min\{1,1/2^{j}\}]^{p}\biggr]^{1/p}$
with $\frac{1}{p}+\frac{1}{q}=1$. Note that from this inequality it follows that
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot p_{\theta,q}(x,y).$ (12)
Furthermore, for given $x$ and $y$, the constant approximating sequences $x_{n}=x$ and $y_{n}=y$ for all natural $n$ are such that for these, the bi-infinite linking sequence $(S_{k}^{x,y})$ satisfies:
$S_{k}^{x,y}=\begin{cases}x,&\text{if $k>0$}\\ y,&\text{if $k\leq 0$}\end{cases}$
Thus, for this bi-infinite sequence, we have
$\sum_{k\in\mathbb{Z}}[2^{-k\theta}\cdot J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y})]^{q}=[J_{M}(1;x,y)]^{q}.$
From this, it follows that
$p_{\theta,q}(x,y)\leq\Gamma_{\theta,q}((J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y}))\leq J_{M}(1;x,y).$ (13)
From inequality 13, we have that $p_{\theta,q}(x,y)<+\infty$ for all $x,y\in X_{0}\cap X_{1}$ and from inequality 12 we have that $p_{\theta,q}(x,y)=0\Longleftrightarrow x=y$. ∎
Lemma 3.2.
Let $\theta\in(0,1)$ and $q\in[1,\infty]$. Given $x,y\in X_{0}\cap X_{1}$, let us define the functional
$P_{\theta,q}(x,y):=\frac{p_{\theta,q}(x,y)+p_{\theta,q}(y,x)}{2}.$ (14)
The functional $P_{\theta,q}:(X_{0}\cap X_{1})\times(X_{0}\cap X_{1})\to[0,\infty)$ satisfies
  1. 1.
    $P_{\theta,q}(x,y)=0\Longleftrightarrow x=y$.
  2. 2.
    $P_{\theta,q}(x,y)=P_{\theta,q}(y,x)$, for all $x,y\in X_{0}\cap X_{1}$.
  3. 3.
    $\min\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}\leq P_{\theta,q}(x,y)\leq\max\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}$, for all $x,y\in X_{0}\cap X_{1}$.
Proof.
The proof is immediate. ∎
Definition 3.3.
Given $\theta\in(0,1)$ and $q\in[1,\infty]$ let us define the functional $\delta_{\theta,q}$ on $X_{0}\cap X_{1}$, given by
$\delta_{\theta,q}(x,y):=\inf\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1}),$ (15)
where the infimum is taken over all finite sequences of points $x_{0}=x,x_{1},x_{2},x_{3},\ \ldots,x_{n}=y\in X_{0}\cap X_{1}$.
Theorem 3.3.
The functional $\delta_{\theta,q}$ as above is a metric on $X_{0}\cap X_{1}$.
Proof.
It is easy to verify that $\delta_{\theta,q}(x,x)=0$. Furthermore, for all $x,y\in X_{0}\cap X_{1}$, taking the sequence $x_{0}=x$ and $x_{1}=y$, we have
$0\leq\delta_{\theta,q}(x,y)\leq P_{\theta,q}(x,y)<+\infty.$
Let now $x,y\in X_{0}\cap X_{1}$ such that $\delta_{\theta,q}(x,y)=0$. From inequality 12, we have
$K_{M}(1;x,y)=K_{M}(1;y,x)\leq M_{\theta,q}\cdot p_{\theta,q}(x,y),$
for all $x,y\in X_{0}\cap X_{1}$. Thus, it is also true that $K_{M}(1;x,y)\leq M_{\theta,q}\cdot p_{\theta,q}(y,x)$. Therefore,
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot\min\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}\leq M_{\theta,q}\cdot P_{\theta,q}(x,y),$ (16)
for any $x,y\in X_{0}\cap X_{1}$. For any linking sequence $x_{0}=x,x_{1},x_{2},\ldots,x_{n}=y$ we have
$K_{M}(1;x,y)$ $\leq\sum_{j=0}^{n-1}K_{M}(1;x_{j},x_{j+1})$
$\leq M_{\theta,q}\cdot\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1}).$
Consequently, we have
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot\delta_{\theta,q}(x,y),\quad x,y\in X_{0}\cap X_{1}.$ (17)
This inequality guarantees the positivity of $\delta_{\theta,q}$. The symmetry of $\delta_{\theta,q}$ comes from the symmetry of $P_{\theta,q}$ and the fact that every linking sequence from $x$ to $y$ in $X_{0}\cap X_{1}$ is also a linking sequence from $y$ to $x$ in $X_{0}\cap X_{1}$. For the triangle inequality, let $x,y,z\in X_{0}\cap X_{1}$ and $\varepsilon>0$. Then, there exist linking sequences $x_{0}=x,x_{1},x_{2},\ldots,x_{n}=z$ from $x$ to $z$ and $y_{0}=z,y_{1},y_{2},\ldots,y_{m}=y$ from $z$ to $y$, such that
$\delta_{\theta,q}(x,z)\leq\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1})<\delta_{\theta,q}(x,z)+\varepsilon$
and
$\delta_{\theta,q}(z,y)\leq\sum_{j=0}^{m-1}P_{\theta,q}(y_{j},j_{j+1})<\delta_{\theta,q}(z,y)+\varepsilon.$
Note that the sequence
$z_{0}=x,z_{1},\ldots,z_{n}=z,z_{n+1}=z,z_{n+2},\ldots,z_{n+m}=y$
is a linking sequence from $x$ to $y$ in $X_{0}\cap X_{1}$ and satisfies
$\delta_{\theta,q}(x,y)$ $\leq\sum_{j=0}^{m+n-1}P_{\theta,q}(z_{j},z_{j+1})$
$=\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1})+\sum_{j=0}^{m-1}P_{\theta,q}(y_{j},y_{j+1})$
$<\delta_{\theta,q}(x,z)+\delta_{\theta,q}(z,y)+2\varepsilon.$
From the arbitrariness of $\varepsilon>0$ it follows
$\delta_{\theta,q}(x,y)\leq\delta_{\theta,q}(x,z)+\delta_{\theta,q}(z,y).$
Therefore, $\delta_{\theta,q}$ is indeed a metric on $X_{0}\cap X_{1}$. ∎
Remark 3.1.
Inequality 17 ensures that the metric space $(X_{0}\cap X_{1},\delta_{\theta,q})$ is continuously included in the metric space $(X_{0}\cup X_{1},K_{M}(1))$, thus satisfying the conditions for relative completion.
Definition 3.4 (The $J_{M}^{\prime}$-interpolated metric space).
The $J_{M}^{\prime}$-interpolated metric space is denoted by $(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$ and is the relative completion of the metric space
$(X_{0}\cap X_{1},\delta_{\theta,q})$
in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. The metric of this space is denoted by $\Delta_{\theta,q}$ and equals $\Delta_{\theta,q}(x,y)=\delta_{\theta,q}(x,y)$, for all $x,y\in X_{0}\cap X_{1}$.
Remark 3.2.
An important fact is that the interpolated spaces above were defined as the completion of the intersection $X_{0}\cap X_{1}$ with a specific metric in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. This ensures that the elements of the interpolated spaces are contained in $X_{0}\cup X_{1}$. However, note that such metrics on the intersection also satisfy the conditions for completion in $(X,d_{X})$ which can be substantially “larger” than $X_{0}\cup X_{1}$, providing many more new elements to the completions. That is why the name ”relative” was chosen for this type of completion. We do not study here the advantages of completing the spaces beyond $X_{0}\cup X_{1}$.
4 Properties of the $J_{M}^{\prime}$-Interpolated Metric Space
Theorem 4.1.
The inclusions $X_{0}\cap X_{1}\subseteq(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}\subseteq X_{0}\cup X_{1}$ are valid and furthermore
  1. 1.
    $\Delta_{\theta,q}(x,y)=\delta_{\theta,q}(x,y)\leq J_{M}(1;x,y),\quad x,y\in X_{0}\cap X_{1}$.
  2. 2.
    $K_{M}(1;x,y)\leq M_{\theta,q}\cdot\Delta_{\theta,q}(x,y),\quad x,y\in(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$.
Proof.
The inclusions are immediate. We will show only the inequalities. Let $x,y\in X_{0}\cap X_{1}$. From inequality 13, we have
$p_{\theta,q}(x,y)\leq J_{M}(1;x,y).$
Since $J_{M}(1;x,y)=J_{M}(1;y,x)$, it holds that
$\max\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}\leq J_{M}(1;x,y).$
Thus, $P_{\theta,q}(x,y)\leq J_{M}(1;x,y)$. For the trivial linking sequence $x_{0}=x$ and $x_{1}=y$, we have
$\delta_{\theta,q}(x,y)\leq P_{\theta,q}(x,y)\leq J_{M}(1;x,y).$ (18)
The second inequality is obtained directly from inequality 17. Therefore, the inclusions are continuous, as we wanted to demonstrate. ∎
Definition 4.1.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces and $T:X\to Y$ an operator. We say that $T$ is a closed operator if
$d_{X}(x_{n},x)\to 0\quad\text{and}\quad d_{Y}(T(x_{n}),y)\to 0]\implies y=T(x).$
In other words, the set
$\text{Gr}(T):=\{(x,T(x))\in X\times Y\}$
is a closed set in $X\times Y$ in the product topology.
Definition 4.2.
Let $(X,d_{X})$, $(y,d_{Y})$ and $T:X\to Y$ an operator. We say that the operator $T$ is a Lipschitz operator if there exists a positive constant $C$ such that
$d_{Y}(T(x),T(y))\leq C\cdot d_{X}(x,y),\quad\forall x,y\in X.$
When $T$ is a Lipschitz operator, we call the number
$\inf_{C>0}\{d_{Y}(T(x),T(y))\leq C\cdot d_{X}(x,y);\quad\forall x,y\in X\}$
the Lipschitz constant of the operator $T$.
Lemma 4.2 (Reindexing and interpolation factor).
Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces and $T:X\to Y$ an operator. Let $(S_{k})_{k\in\mathbb{Z}}$ be a linking sequence from $x$ to $y$ in $X$ (which for simplicity of notation we will omit the superscript $x,y$) and let
$a_{k}:=2^{-k\theta}J_{M}(2^{k}\lambda;\,T(S_{k}),T(S_{k+1}):Y),\qquad b_{k}:=2^{-k\theta}J_{M}(2^{k};\,S_{k},S_{k+1}:X),$
where $\lambda>0$ and $\theta\in(0,1)$. Suppose that for all $k\in\mathbb{Z}$ it holds that
$J_{M}(2^{k};S_{k},S_{k+1}:X)\;\geq\;\omega_{0}^{-1}\,J_{Y}(2^{k}\lambda;\,T(S_{k}),T(S_{k+1}):Y),\quad\omega_{0}>0.$ (H)
For $\omega_{1}>0$ define $\lambda=\dfrac{\omega_{0}}{\omega_{1}}$. Then:
  1. (i)
    If $\lambda=2^{r}$ for some $r\in\mathbb{Z}$, then, for all $q\in[1,\ infty]$,
    $\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\;\big\|(b_{k})_{k}\big\|_{\ell^{q}}.$
  2. (ii)
    If $\lambda>0$ is arbitrary, then, for all $q\in[1,\infty]$,
    $\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq C(\theta)\,\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\;\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
    where $C(\theta)\leq 2^{\theta}$ is a constant depending only on $\theta$.
Proof.
We start by observing that hypothesis (H) is equivalent, in terms of the terms $a_{k},b_{k}$, to
$b_{k}\;\geq\;\omega_{0}^{-1}a_{k},\qquad\text{that is}\qquad a_{k}\leq\omega_{0}\,b_{k},\qquad\forall k\in\mathbb{Z}.$
If $\lambda=2^{r}$, $r\in\mathbb{Z}$: Writing $\lambda=2^{r}$ and making the index change $m=k+r$ we obtain, for each $k$,
$a_{k}=2^{-k\theta}J_{M}(2^{k}\lambda;\,T(S_{k}),T(S_{k+1}):Y)=2^{-k\theta}J_{M}(2^{k+r};\,T(S_{k}),T(S_{k+1}):Y).$
Rewriting in terms of the index $m$ (with $m=k+r$) we have the identity
$a_{k}$ $=2^{r\theta}\,2^{-(k+r)\theta}J_{M}(2^{k+r};\,T(S_{k}),T(S_{k+1}):Y)$
$=\lambda^{\theta}\,2^{-m\theta}J_{M}(2^{m};\,T(S_{m-r}),T(S_{m-r+1}):Y).$
By hypothesis (H) applied to the index mrm-r we have
$2^{-m\theta}J_{M}(2^{m};\,T(S_{m-r}),T(S_{m-r+1}):Y)\leq$
$\omega_{0}\,2^{-(m-r)\theta}J_{M}(2^{m-r};\,S_{m-r},S_{m-r+1}:X).$
Combining the two expressions and using reindexing (the transformation $k\mapsto m=k+r$ is an index shift), we obtain
$2^{-m\theta}J_{M}(2^{m};\,T(S_{m}),T(S_{m+1}):Y)\leq$
$\omega_{0}\,\lambda^{-\theta}\,2^{-(m-r)\theta}J_{M}(2^{m-r};\,S_{m-r},S_{m-r+1}:X).$
Taking the $\ell^{q}$ norm of both sides and remembering that the index shift is isometric in $\ell^{q}$ (that is, $\|(u_{m-r})_{m}\|_{\ell^{q}}=\|(u_{k})_{k}\|_{\ell^{q}}$), it results in
$\big\|(2^{-m\theta}J_{M}(2^{m};T(S_{m}),T(S_{m+1});Y))_{m}\big\|_{\ell^{q}}\leq$
$\omega_{0}\,\lambda^{-\theta}\,\big\|(2^{-k\theta}J_{X}(2^{k};S_{k},S_{k+1}:X))_{k}\big\|_{\ell^{q}}.$
But the left-hand side is exactly $\|(a_{k})_{k}\|_{\ell^{q}}$ (after reindexing), and the right-hand side is $\|(b_{k})_{k}\|_{\ell^{q}}$. Substituting $\lambda=\dfrac{\omega_{0}}{\omega_{1}}$ we obtain
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}\Big(\frac{\omega_{0}}{\omega_{1}}\Big)^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}}=\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
which proves (i).
Arbitrary Case $\lambda>0$: Choose $r=\lfloor\log_{2}\lambda\rfloor\in\mathbb{Z}$ such that $2^{r}\leq\lambda<2^{r+1}$. For all $k$ we have the two elementary estimates
$J_{M}(2^{k+r};\cdot,\cdot:Y)\;\leq\;J_{M}(2^{k}\lambda;\cdot,\cdot:Y)\;\leq\;J_{Y}(2^{k+r+1};\cdot,\cdot:Y),$
because the function $s\mapsto J_{M}(2^{k}s;\cdot,\cdot:Y)$ is non-decreasing in $s>0$ (remembering that here only the multiplicative parameter within the second argument of $J_{M}$ varies). Multiplying by $2^{-k\theta}$ and reindexing as before, we obtain
$2^{r\theta}\,2^{-(k+r)\theta}J_{M}(2^{k+r};\cdot,\cdot:Y)$ $\leq 2^{-k\theta}J_{M}(2^{k}\lambda;\cdot,\cdot:Y)$
$\leq 2^{(r+1)\theta}\,2^{-(k+r+1)\theta}J_{M}(2^{k+r+1};\cdot,\cdot:Y).$
Applying inequality (H) (in the form used in (i)) to the appropriate indices and taking the $\ell^{q}$ norm, we arrive at inequalities of the type
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}\,2^{\theta}\,\lambda^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}}\quad\text{or}\quad\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}\,\lambda^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
depending on whether we use the majorant or the minorant above. In both cases we can write compactly
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq C(\theta)\,\omega_{0}\,\lambda^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},\qquad C(\theta)\leq 2^{\theta}.$
Substituting $\lambda=\dfrac{\omega_{0}}{\omega_{1}}$ results in
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq C(\theta)\,\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
as we wanted to demonstrate. ∎
Theorem 4.3.
Let $(X_{0},d_{X_{0}})$ and $(X_{1},d_{X_{1}})$ be a pair of compatible metric spaces with respect to the metric space $(X,d_{X})$ and $(Y_{0},d_{Y_{0}})$ and $(Y_{1},d_{Y_{1}})$ a pair of compatible metric spaces with respect to the metric space $(Y,d_{Y})$. If $(Y,d_{Y})$ is complete and $T:X\to Y$ is a closed operator that satisfies:
  1. 1.
    $T_{0}:=T|_{X_{0}}:X_{0}\to Y_{0}$ is Lipschitzian with Lipschitz constant $\omega_{0}>0$
  2. 2.
    $T_{1}:=T|_{X_{1}}:X_{1}\to Y_{1}$ is Lipschitzian with Lipschitz constant $\omega_{1}>0$
Then, $T_{\theta,q}:=T|_{(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}}:(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}\to(\overrightarrow{Y}_{\theta,q}^{J_{M}})^{\prime}$ is Lipschitzian with Lipschitz constant not exceeding $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}>0$.
Proof.
Let $x,y\in X_{0}\cap X_{1}$ and $(x_{n})$ and $(y_{n})$ be two approximating sequences of $x$ and $y$ in $X_{0}\cap X_{1}$. Let us define the bi-infinite linking sequence $(S_{k}^{x,y})$ that links $x$ to $y$ in $X_{0}\cap X_{1}$ as usual through these. Since $(x_{n})$ is Cauchy in $(X_{0}\cap X_{1},J_{M}(1))$, it follows that $(x_{n})$ is Cauchy in $(X_{0},d_{X_{0}})$ and in $(X_{1},d_{X_{1}})$, because
$J_{M}(1;x_{m},x_{n}:X)$ $:=\max\{d_{X_{0}}(x_{m},x_{n}),d_{X_{1}}(x_{m},x_{n})\}$
$\geq d_{X_{i}}(x_{m},x_{n}),\quad i=0,1.$
Since $(x_{n})$ is Cauchy in $(X_{i},d_{X_{i}})$, $i=0,1$, due to the fact that
$d_{Y_{i}}(T_{i}(x_{m}),T_{i}(x_{n}))\leq\omega_{i}\cdot d_{X_{i}}(x_{m},x_{n}),\quad i=0,1,$
we have that $(T(x_{n}))$ is a Cauchy sequence both in $(Y_{0},d_{Y_{0}})$ and in $(Y_{1},d_{Y_{1}})$. Since by hypothesis there exists $C_{i}>0$ such that $d_{Y}(T(x_{m}),T(x_{n}))\leq C_{i}\cdot d_{Y_{i}}(T(x_{m}),T(x_{n}))$, for $i=0,1$, we have that $(T(x_{n}))$ is a Cauchy sequence in $(Y,d_{Y})$, which is complete, by hypothesis. It follows that there exists $a\in Y$ such that $d_{Y}(T(x_{n}),a)\to 0$. By the closedness property of the operator $T$, we have $a=T(x)$. The same reasoning can be applied to the sequence $(y_{n})$ which will provide $b\in Y$ such that $T(y)=b$. Note that if $(T(S_{k}^{x,y}))$ is a bi-infinite linking sequence from $T(y)$ to $T(x)$ in $Y_{0}\cap Y_{1}$. Thus, since
$J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y}:X)=$
$\max\{d_{X_{0}}(S_{k}^{x,y},S_{k+1}^{x,y});2^{k}d_{X_{1}}(S_{k}^{x,y},S_{k+1}^{x,y})\}\geq$
$\max\{\omega_{0}^{-1}d_{Y_{0}}(T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)}),2^{k}\omega_{1}^{-1}d_{Y_{1}}(T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)})\}=$
$\omega_{0}^{-1}\cdot J_{M}(2^{k}\omega_{0}/\omega_{1};T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)}):Y)\geq$
$2^{-\theta}\cdot\omega_{0}^{\theta-1}\cdot\omega_{1}^{-\theta}\cdot J_{M}(2^{k};T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)}):Y),$
where in the last inequality we used the previous Lemma 4.2. From this inequality, we have that
$p_{\theta,q;Y}(T(x),T(y))\leq 2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}\cdot p_{\theta,q;X}(x,y).$
From this inequality, it is immediate that
$\Delta_{\theta,q;Y}(T(x),T(y))\leq 2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$ (19)
And this concludes the proof. ∎
Remark 4.1.
If we use the result that
$J_{M}(t;\cdot,\cdot)\leq\max\{1,t/s\}\cdot J_{M}(s;\cdot,\cdot)$
we can eliminate the dependence on $\theta$ from the upper bound of the Lipschitz constant of the operator $T$ in the interpolated space. In this situation, setting $t=2^{k}$ and $s=2^{k}\omega_{0}/\omega_{1}$, we have
$\max\{\omega_{0},\omega_{1}\}\cdot J_{M}(2^{k};S_{k},S_{k+1}:X)\geq J_{M}(2^{k};T(S_{k}),T(S_{k+1}):Y),$
resulting in
$\Delta_{\theta,q;Y}(T(x),T(y))\leq\max\{\omega_{0},\omega_{1}\}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}.$ (20)
We choose to display the possible dependence on $\theta$ because it is a constant more characteristic of the language in Interpolation Theory and expresses some type of convexity.
Corollary 4.4.
In Theorem 4.3 consider $X=Y$ and $d_{X}=d_{Y}$ a metric space complete. If $\omega_{i}<1$ for $i=0,1$, and $(X_{0},d_{X_{0}})$ and $(X_{1},d_{X_{1}})$ are both complete then there exists a unique fixed point of $T$ in $X_{0}\cup X_{1}$, even if $T$ is not necessarily Lipschitzian nor continuous in $(X_{0}\cup X_{1},K_{M}(1))$.
Proof.
Since $\max\{\omega_{0},\omega_{1}\}<1$, it follows from inequality 20 that $T_{\theta,q}$ (the operator $T$ restricted to the interpolated space) has a unique fixed point $x^{\prime}$, by the Banach Fixed Point Theorem. Such a fixed point necessarily belongs to $X_{0}\cup X_{1}$. Thus, $x^{\prime}\in X_{0}$ or $x^{\prime}\in X_{1}$. Since by hypothesis the operator $T$ is a strong contraction in $X_{0}$ and in $X_{1}$, unique fixed points $x_{0}\in X_{0}$ and $x_{1}\in X_{1}$ exist. Therefore, we must have $x^{\prime}=x_{0}=x_{1}$ and no other distinct point can be fixed by $T$ in $X_{0}\cup X_{1}$. ∎
Remark 4.2.
Throughout the proof we used the same letter $T$ to indicate different operators, since these operators were only restrictions of the same $T$ to spaces that were clear in the context through other symbols.
Remark 4.3.
If we impose the condition of continuity on $T:X\to Y$ we can waive the completeness of $(Y,d_{Y})$.
Definition 4.3.
Let $T:X\to Y$ be an operator between metric spaces. We will say that $T$ is a compact operator if for every sequence $(x_{n})$ of points in $X$ that is bounded, there exists a subsequence $(x_{n_{k}})$ of $(x_{n})$ such that $(T(x_{n_{k}}))$ is convergent.
Proposition 4.5.
If $X_{0}=X_{1}=X$ and $d_{0}=d_{1}=d_{X}$. Given $\theta\in(0,1)$ and $q\in[1,\infty]$, we have
$(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}=X$
and the metric $\Delta_{\theta,q}$ is equivalent to the metric $d_{X}$.
Proof.
For $t>0$ and $u,v\in X$,
$J_{M}(t;u,v)=\max\{d_{X}(u,v),\,t\,d_{X}(u,v)\}=\begin{cases}d_{X}(u,v),&0<t\leq 1,\\[4.0pt] t\,d_{X}(u,v),&t\geq 1.\end{cases}$
In particular, for $t=1$ we have $J_{M}(1;u,v)=d_{X}(u,v)$, and the metric $K_{M}(1;\cdot,\cdot)$ defined by infimum over linking sequences coincides with the original metric $d_{X}$ (since the cost per step is always $h_{1}(\cdot,\cdot)=d_{X}(\cdot,\cdot)$). By inequalities (12) and (13)):
$K_{M}(1;u,v)\leq M_{\theta,q}\,p_{\theta,q}(u,v),$ (21)
and for constant sequences,
$p_{\theta,q}(u,v)\leq J_{M}(1;u,v).$ (22)
In the trivial case, $K_{M}(1;u,v)=d_{X}(u,v)$ and $J_{M}(1;u,v)=d_{X}(u,v)$. Substituting these equalities into (21) and (22) we obtain, for all $u,v\in X$,
$d_{X}(u,v)\leq M_{\theta,q}\,p_{\theta,q}(u,v)\qquad\text{and}\qquad p_{\theta,q}(u,v)\leq d_{X}(u,v).$
Rewriting,
$\frac{1}{M_{\theta,q}}\,d_{X}(u,v)\;\leq\;p_{\theta,q}(u,v)\;\leq\;d_{X}(u,v).$ (23)
The inequalities (23) show that $p_{\theta,q}$ is equivalent to the metric $d_{X}$ (with explicit constants $1/M_{\theta,q}$ and $1$). We have
$P_{\theta,q}(u,v)=\tfrac{1}{2}\bigl(p_{\theta,q}(u,v)+p_{\theta,q}(v,u)\bigr),$
and
$\delta_{\theta,q}(x,y)=\inf\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1}),$
where the infimum runs over all linking sequences $x_{0}=x,\dots,x_{n}=y$ in $X_{0}\cap X_{1}=X$. Since $p_{\theta,q}$ satisfies (23), the same inequalities hold for $P_{\theta,q}$. More precisely, using (23) and the definition of $P_{\theta,q}$, we obtain
$\frac{1}{M_{\theta,q}}\,d_{X}(u,v)\leq P_{\theta,q}(u,v)\leq d_{X}(u,v),\qquad\forall u,v\in X.$
Now, by the definition of $\delta_{\theta,q}$, we have
$\frac{1}{M_{\theta,q}}\,d_{X}(x,y)\leq\delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ inX.$
Therefore $\delta_{\theta,q}$ is equivalent to $d_{X}$. Since $(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$ was defined as the relative completion of $(X_{0}\cap X_{1},\delta_{\theta,q})$ in $(X_{0}\cup X_{1},K_{M}(1))$ and here $X_{0}\cap X_{1}=X_{0}\cup X_{1}=X$, it holds that
$(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}=X.$
Since $\Delta_{\theta,q}=\delta_{\theta,q}$ on $X_{0}\cap X_{1}=X$, we have
$\frac{1}{M_{\theta,q}}\,d_{X}(x,y)\leq\Delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ \in X$
as we wanted to demonstrate. ∎
Corollary 4.6.
Let $(X_{0},X_{1})_{X}$ be a pair of metric spaces, $\theta\in(0,1)$, $q\in[1,\infty]$ and $(F,d_{F})$ a complete metric space. If $T$ is an operator that satisfies
  • (i)
    $T:X\rightarrow F$ is closed;
  • (ii)
    $T:X_{0}\rightarrow F$ is a compact Lipschitzian operator;
  • (iii)
    $T:X_{1}\rightarrow F$ is Lipschitzian;
  • (iv)
    Every sequence bounded in $X_{0}\cap X_{1}$ with respect to the interpolated metric is also bounded with respect to the metric $d_{X_{0}}$,
then $T_{\theta,q}:=T|_{(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}}:(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}\to F$ is Lipschitzian and compact.
Proof.
The result is immediate from the definitions, proposition 4.5 and theorem 4.3. ∎
Remark 4.4.
The metric interpolation developed in this work can interact with non-linear structures relevant in analysis and PDEs. A natural example is the space of probabilities $\mathcal{P}_{2}(\Omega)$ equipped with the Wasserstein distance $W_{2}$, which is a complete and geodesic metric space, but non-linear. In many evolution problems — such as gradient flows associated with energy functionals in the sense of Jordan–Kinderlehrer–Otto — regularity on different metric scales plays a crucial role. The simultaneous existence of two metric controls (one more “fine”, the other more “coarse”) suggests the possibility of applying metric interpolation constructions to obtain intermediate estimates of stability or temporal continuity.
Although such applications exceed the scope of this article, the abstract structure developed here provides a framework that can be used to investigate gradient flows, metric compactifications, and regularization properties in Wasserstein-type spaces.