### v2

# A new interpolation method for metric spaces based on bi-infinite sequences: The $R$-Method

Abstract
We introduce a new interpolation method for metric spaces, termed the $R$-method, based on bi-infinite linking sequences. Although the construction is inspired by the classical metric functional $J_{M}$, the resulting interpolated space is generated by a distinct object that behaves as a multiscale energy functional. This functional measures the minimal discrete action required to connect two points through $\mathbb{Z}$-indexed sequences, leading to a new intrinsic metric on $X_{0}\cap X_{1}$.
The associated interpolated space is obtained as the relative completion of this metric inside $X_{0}\cup X_{1}$ and is genuinely different from those produced by the $J_{M}$- and $K_{M}$-methods. A fundamental structural property of the $R$-method is that the resulting space embeds continuously into the corresponding $K_{M}$-interpolated space, situating the construction naturally within the existing theory of metric interpolation.
When the method is restricted to a normed setting, the $R$-method induces a genuine interpolation functor. In this framework, it preserves the Lipschitz property of operators with closed graphs, even in the absence of linearity, thereby extending the classical scope of interpolation theory, which is traditionally confined to linear continuous operators. As a consequence, standard compactness properties are also preserved under mild assumptions.
The $R$-method thus provides a new interpolation framework whose foundations rely exclusively on intrinsic metric properties and the summability of discrete orbits, bridging metric interpolation, nonlinear analysis, and classical interpolation theory.
Keywords:
Metric Interpolation. Lipschitz Operators. Nonlinear Analysis. J-method. K-method. Functional Analysis. Interpolation Theory.
Classification codes:
46B70, 54E35, 47H09.
Introduction
The classical theory of interpolation, initiated in the 1960s through the foundational work of Lions and Peetre [10] and later systematized in texts such as [3, 4], is intrinsically rooted in linear and normed structures. Interpolation functors were designed to construct intermediate Banach spaces between two given endpoints, providing powerful tools for the analysis of linear operators, partial differential equations, harmonic analysis, and geometric functional analysis. For decades, the theory evolved almost exclusively within linear frameworks, with its principal constructions relying on convexity, homogeneity, and duality.
Only recently has the need for interpolation tools beyond the linear setting become apparent. Several modern areas, including optimal transport, metric geometry, gradient flows in Wasserstein spaces, and nonlinear analysis on general metric measure spaces, are naturally formulated in purely metric terms and lack vector space structure. Nevertheless, many of these contexts exhibit intermediate regularity phenomena that are conceptually analogous to interpolation. This has motivated the search for interpolation frameworks that operate directly at the level of metric spaces, without recourse to linear or normed structures. Despite progress in nonlinear functional analysis and metric geometry, a systematic and flexible interpolation theory for arbitrary metric spaces has remained largely undeveloped.
In [13], we initiated such a program by introducing two metric-space analogues of Peetre’s classical $K$- and $J$-methods, denoted respectively by the $K_{M}$- and $J_{M}$-methods. These constructions reinterpret the classical ideas entirely in metric terms. Given a compatible pair of metric spaces $(X_{0},X_{1})_{X}$, the $K_{M}$-method is based on admissible finite linking sequences in $X_{0}\cup X_{1}$, penalizing each step through a functional of the form
$K_{M}(t;x,y)=\inf\sum h_{t}(x_{k},x_{k+1}),$
which mirrors the behavior of the classical $K$-functional. In contrast, the $J_{M}$-method acts on the intersection $X_{0}\cap X_{1}$ and is governed by the metric
$J_{M}(t;x,y)=\max\{d_{0}(x,y),\,t\,d_{1}(x,y)\},$
the natural metric counterpart of Peetre’s $J$-functional.
Both methods generate interpolated metric spaces via relative completion: one first endows $X_{0}\cap X_{1}$ with a suitable multiscale metric functional and then completes it inside the ambient space $(X_{0}\cup X_{1},K_{M}(1))$. These constructions preserve Lipschitz continuity of operators, exhibit functorial behavior analogous to that of classical interpolation methods, and apply naturally in purely metric settings, including Wasserstein spaces [1, 15, 16]. Their introduction represents one of the first systematic attempts to extend interpolation theory beyond the linear world.
The present article introduces a third interpolation method, which is structurally distinct from both the $K_{M}$- and $J_{M}$-methods. Although the construction still involves the classical functional $J_{M}$, the underlying mechanism is fundamentally different. Rather than relying on finite admissible chains or single-step comparisons, the new method is built upon bi-infinite linking sequences
$(S_{k})_{k\in\mathbb{Z}}\subset X_{0}\cap X_{1},$
which simultaneously approximate a point $x$ as $k\to+\infty$ and a point $y$ as $k\to$-$\infty$. Along such sequences, we evaluate a multiscale functional of the form
$\Gamma_{\theta,q}\bigl(J_{M}(2^{k};S_{k},S_{k+1})\bigr),$
measuring the minimal discrete action required to connect $x$ and $y$ across all dyadic scales. Taking the infimum over all admissible approximating sequences yields a pre-metric $p_{\theta,q}$; symmetrization and path-length completion then produce a genuine metric $\delta_{\theta,q}$ on $X_{0}\cap X_{1}$.
The resulting interpolated space is defined as the relative completion of $(X_{0}\cap X_{1},\delta_{\theta,q})$ inside $(X_{0}\cup X_{1},K_{M}(1))$. This construction yields a new metric interpolation space, denoted by $R_{\theta,q}^{\overrightarrow{X}}$, whose geometry differs essentially from that of the previously studied $K_{M}$- and $J_{M}$-interpolated spaces. The use of bi-infinite approximating sequences introduces a variational and multiscale structure, capturing information distributed across all dyadic scales. In this respect, the method is closer in spirit to energetic approaches common in modern metric analysis, while remaining firmly rooted in interpolation theory.
From a broader perspective, the development of interpolation theory for metric spaces remains in its early stages, and the availability of multiple mechanisms is essential. Different analytic and geometric problems may require different ways of measuring intermediate regularity or oscillation. The $R$-method enriches the existing toolkit by providing a sequence-based, variational, and genuinely multiscale interpolation mechanism, which may be advantageous in nonlinear settings where convergence of sequences plays a central role, such as gradient flows, geometric evolution problems, and approximation schemes in metric geometry.
The purpose of this article is to present the full construction of the $R$-method, establish its fundamental metric properties, and show that it is compatible with the interpolation of Lipschitz and compact operators. While the method is intrinsically metric, we also show that, when restricted to normed spaces, it induces a genuine interpolation functor, extending the classical scope of interpolation beyond linear operators. In doing so, the article lays the foundations for a new and distinct interpolation mechanism within the emerging theory of metric-space interpolation.
1 Preliminaries
All results and definitions in this section are found in the work [see [13]]. Thus, we decided to omit the proofs of basic lemmas of the theory that fall outside the scope of this text. The interested reader can verify the proofs in the reference above.
1.1 Relative Completion of Metric Spaces
Definition 1.1.
Let $(A,d_{A})$ and $(B,d_{B})$ be two metric spaces such that $A\subseteq B$ and there exists $c>0$ such that
$d_{B}(x,y)\leq c\cdot d_{A}(x,y),\quad x,y\in A.$
We define the relative completion of $(A,d_{A})$ in $(B,d_{B})$, denoted by $A_{B}$, as the set of all the $x\in B$ such that there exists a Cauchy sequence $(x_{n})$ in $(A,d_{A})$ satisfying $d_{B}(x_{n},x)\to 0$.
A sequence $(x_{n})$ as above is called an approximating sequence of $x$ in $(A,d_{A})$.
There are some intrinsic properties of the relative completion of metric spaces that we list below:
  1. 1.
    This completion can be metrized with a metric $d_{A,B}$ that satisfies
    $d_{B}(x,y)\leq c\cdot d_{A,B}(x,y)\leq d_{A}(x,y),\quad x,y\in A$
    and
    $d_{B}(x,y)\leq c\cdot d_{A,B}(x,y),\quad x,y\in A_{B}.$
    Thus, $A\subseteq A_{B}\subseteq B$ with continuous inclusions.
  2. 2.
    The set $A$ is dense in $(A_{B},d_{A,B})$.
  3. 3.
    If $(B,d_{B})$ is complete, then $(A_{B},d_{A,B})$ is also complete.
  4. 4.
    The set $A$ is dense in $(A_{B},d_{A,B})$.
  5. 5.
    The metrics $d_{A,B}$ and $d_{A}$ coincide on $A$.
1.2 Pairs of Compatible Metric Spaces
Definition 1.2.
Let $X_{0}$ and $X_{1}$ be two sets such that $X_{0}\cap X_{1}\neq\emptyset$ and both are subsets of a set $X$. Furthermore, assume that $d_{0},d_{1}$ and $d_{X}$ are metrics on the sets $X_{0}$, $X_{1}$ and $X$, respectively, such that for convenient positive constants $C_{0}$ and $C_{1}$ we have
$d_{X}(x,y)\leq C_{i}\cdot d_{i}(x,y),\quad\forall\,x,y\in X_{i},\quad i\in\{0,1\}.$ (1)
We call this object a pair of compatible metric spaces and denote it by $\overrightarrow{X}$ or by $(X_{0},X_{1})_{X}$.
Let $(X_{0},X_{1})_{X}$ and $t>0$. For $(x,y)\in X_{0}^{2}\cup X_{1}^{2}$, define the functional hop, denoted by $h_{t}$ which will be given by
$h_{t}(x,y)=\left\{\begin{array}[c]{l}\min\{d_{0}(x,y),t\cdot d_{1}(x,y)\},\text{ if }x,y\in X_{0}\cap X_{1};\\ d_{0}(x,y),\text{ if}\,\,\{x,y\}\subseteq X_{0}\setminus X_{1};\\ t\cdot d_{1}(x,y),\text{ if}\,\,\{x,y\}\subseteq X_{1}\setminus X_{0}.\end{array}\right$.
Definition 1.3.
Given $x,y\in X_{0}\cup X_{1}$, with $X_{0}\cap X_{1}\neq\emptyset$, an admissible linking sequence from $x$ to $y$ in $X_{0}\cup X_{1}$ is any finite sequence of points $x_{0},x_{1},\ldots,x_{n}\in X_{0}\cup X_{1}$ that satisfies:
  • $x_{0}=x$ and $x_{n}=y$;
  • For each $k\in\{0,1,\ldots,n-1\}$, $(x_{k},x_{k+1})\in X_{0}^{2}\ cupX_{1}^{2}$ holds.
Let us denote an admissible linking sequence $(x,x_{1},x_{2},\ldots,x_{n-1},y)$ from $x$ to $y$ in $X_{0}\cup X_{1}$ by $(x_{n}^{x,y})_{adm}$. When all $n$ points of the sequence belong to the same set $Z$, we will say that it is a linking sequence from $x$ to $y$ in $Z$ and denote it only by $(x_{n}^{x,y})$. The set of all linking sequences from $x$ to $y$ in $Z$ is denoted by $x,y]_{Z}$.
Remark 1.
Note that every admissible linking sequence is a linking sequence in the union $X_{0}\cup X_{1}$, but not every linking sequence in $X_{0}\cup X_{1}$ is admissible.
1.3 The $K_{M}$ Functional and its properties
Definition 1.4 (The $K_{M}$-Metric Functional).
Let $(X_{0},X_{1})_{X}$ be a pair of compatible metric spaces and $t>0$. Given $x,y\in X_{0}\cup X_{1}$, we define the $K_{M}$-metric functional by
$K_{M}(t;x,y):=\inf\,\sum_{k=0}^{n-1}h_{t}(x_{k},x_{k+1}),$ (2)
where the infimum is taken over all admissible linking sequences from $x$ to $y$ in the set $X_{0}\cup X_{1}$.
From now on, unless otherwise stated and when there is no possibility of confusion due to notation, $(X_{0},d_{0})$ and $(X_{1},d_{1})$ will be considered a pair of compatible metric spaces with respect to the space $(X,d_{X})$, that is, they will represent $(X_{0},X_{1})_{X}$.
Lemma 1.
[…]
Lemma 2.
If $a,b\in(0,\infty)$ and $x,y\in X_{0}\cup X_{1}$, then
$K_{M}(a;x,y)\leq\max\biggr\{1,\frac{a}{b}\biggr\}\cdot K_{M}(b;x,y)$ (3)
1.4 The $J_{M}$ Functional
Lemma 3.
If $x,y\in X_{0}\cap X_{1}$ and $t>0$, defining
$J_{M}(t;x,y):=\max\{d_{0}(x,y),t\cdot d_{1}(x,y)\},$
[…]
Lemma 4.
If $a,b\in(0,\infty)$ and $x,y\in X_{0}\cap X_{1}$, then
$J_{M}(a;x,y)\leq\max\biggr\{1,\frac{a}{b}\biggr\}\cdot J_{M}(b;x,y),\quad x,y\in X_{0}\cap X_{1};$ (4)
Lemma 5.
If $x,y\in X_{0}\cap X_{1}$ and $a,b\in(0,\infty)$, then
$K_{M}(a;x,y)\leq\min\biggr\{1,\frac{a}{b}\biggr\}\cdot J_{M}(b;x,y).$ (5)
2 The $K_{M}$-Metric Interpolation Space
In this section we will define and highlight the main properties of the interpolated metric space.
Definition 2.1.
Let $\theta\in(0,1)$ and $q\in[1,\infty]$. Given a sequence $(x_{k})_{k\in\ \mathbb{Z}}$ of real numbers, we define the functional $\Gamma_{\theta,q}$ by
$\Gamma_{\theta,q}((x_{k})_{k\in\mathbb{Z}}):=\begin{cases}\biggr[\sum_{k\in\mathbb{Z}}[2^{-k\theta}|x_{k}|]^{q}\biggr]^{1/q},&q<\infty\\ \sup_{k\in\mathbb{Z}}2^{-k\theta}|x_{k}|,&q=\infty.\end{cases}$ (6)
Lemma 6.
The functional
$\beta_{\theta,q}(x,y):=\Gamma_{\theta,q}((K_{M}(2^{k};x,y)_{k\in\mathbb{Z}})$ (7)
is a metric on the set $X_{0}\cap X_{1}$.
We will refer, for simplicity, to the metrics $J_{M}(t;\cdot,\cdot)$ and $K_{M}(t;\cdot,\cdot)$ as $J_{M}(t)$ and $K_{M}(t)$, respectively.
Definition 2.2 (The $K_{M}$-interpolated metric space).
The interpolated metric space by the classical $K_{M}$ method is denoted by $\overrightarrow{X}_{\theta,q}^{K_{M}}$ and is the relative completion of the metric space
$(X_{0}\cap X_{1},\beta_{\theta,q})$
in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. The metric of this space is denoted by $\mathcal{D}_{\theta,q}$ and equals $\mathcal{D}_{\theta,q}(x,y)=\beta_{\theta,q}(x,y)$, for all $x,y\in X_{0}\cap X_{1}$.
Lemma 7.
Given $\theta\in(0,1)$ and $q\geq 1$, we have
$\beta_{\theta,q}(x,y)\leq M_{\theta,q}\cdot J_{M}(1;x,y),\quad x,y\in X_{0}\cap X_{1}.$ (8)
And furthermore,
$M_{\theta,q}\cdot d_{X}(x,y)\leq\mathcal{D}_{\theta,q}(x,y),\quad x,y\in\overrightarrow{X}_{\theta,q}^{K_{M}}.$ (9)
where
$M_{\theta,q}=\displaystyle\Gamma_{\theta,q}(\min\{1,2^{k}\}_{k\in\mathbb{Z}})>0.$
3 The $J_{M}^{\prime}$-Interpolated Metric Space
Definition 3.1 (Bi-infinite Sequences).
A sequence is said to be bi-infinite when it is indexed in the set of integers. A bi-infinite sequence $(z_{k})_{k\in\mathbb{Z}}$ is said to converge to the point $z$ in a metric space $(M,d)$ if
$\lim_{n\to\infty}d(z_{n},z)=0\quad\text{and}\quad\lim_{n\to-\infty}d(z_{n},z)=0.$
We know that $X_{0}\cap X_{1}\subseteq X_{0}\cup X_{1}$ and that $K_{M}(1)$ is a metric on $X_{0}\cup X_{1}$. We will metrize the set $X_{0}\cap X_{1}$ with a metric satisfying the relative completion condition and then complete this space equipped with this metric, with respect to the space $(X_{0}\cup X_{1},K_{M}(1))$.
Given $x,y\in X_{0}\cap X_{1}$ and $(x_{n})$ and $(y_{n})$ approximating sequences of $x$ and $y$ in $(X_{0}\cap X_{1},J_{M}(1))$, respectively, let us define a new bi-infinite sequence, denoted by $(S_{k}^{x,y})_{k\in\mathbb{Z}}$, which depends on these approximating sequences and whose general term is given by:
$S_{k}^{x,y}:=\begin{cases}x_{k},&\text{if $k\geq 0$}\\ y_{-k},&\text{if $k<0$}\end{cases}$
Such a sequence will be called a bi-infinite linking sequence from $x$ to $y$ in $(X_{0}\cap X_{1},J_{M}(1))$. On $X_{0}\cap X_{1}$ let us define the functional $p_{\theta,q}$ given by
$p_{\theta,q}(x,y):=\inf\Gamma_{\theta,q}((J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y})),$ (10)
where the infimum is taken over all approximating sequences of $x$ and $y$.
Definition 3.2 (Separator Functional).
Let $M\neq\emptyset$ and $p:M\times M\to[0,\infty)$. We say that $p$ is a separator functional of points in $M$ and
$p(x,y)=0\Longleftrightarrow x=y.$
Theorem 3.1.
The functional $p_{\theta,q}$ is a separator on $X_{0}\cap X_{1}$.
Proof.
Let $x,y\in X_{0}\cap X_{1}$ and $(x_{n})$ and $(y_{n})$ be two arbitrary approximating sequences of $x$ and $y$ in $(X_{0}\cap X_{1},J_{M}(1))$ and let $(S_{j}^{x,y})_{j\in\mathbb{Z}}$ be the bi-infinite linking sequence from $x$ to $y$ in $(X_{0}\cap X_{1},J_{M}(1))$. Fixing $N\in\mathbb{N}$, we have
$K_{M}(1;S_{N}^{x,y},S_{-N}^{x,y})\leq\sum_{j=-N}^{N-1}K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y}).$
Thus, taking the $\lim\sup_{N}$, we have
$K_{M}(1;x,y)\leq\sum_{j\in\mathbb{Z}}K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y}).$
Now, using Lemma 5, with $a=1$ and $b=2^{j}$ in each term $K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y})$, we have
$K_{M}(1;x,y)\leq\sum_{j\in\mathbb{Z}}K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y})\leq\sum_{j\in\mathbb{Z}}\min\{1,1/2^{j}\}\cdot J_{M}(2^{j};S_{j}^{x,y},S_{j+1}^{x,y}).$
We can also write, for all $\theta\in(0,1)$:
$K_{M}(1;x,y)\leq\sum_{j\in\mathbb{Z}}[2^{j\theta}\cdot\min\{1,1/2^{j}\}]\cdot[2^{-j\theta}\cdot J_{M}(2^{j};S_{j}^{x,y},S_{j+1}^{x,y})].$
Using the Hölder Inequality in the last sum, we have
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot\Gamma_{\theta,q}((J_{M}(2^{j};S_{j}^{x,y},S_{j+1}^{x,y})_{j\in\mathbb{Z}}),$ (11)
where
$M_{\theta,q}:=\biggr[\sum_{j\in\mathbb{Z}}[2^{j\theta}\cdot\min\{1,1/2^{j}\}]^{p}\biggr]^{1/p}$
with $\frac{1}{p}+\frac{1}{q}=1$. Note that from this inequality it follows that
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot p_{\theta,q}(x,y).$ (12)
Furthermore, for given $x$ and $y$, the constant approximating sequences $x_{n}=x$ and $y_{n}=y$ for all natural $n$ are such that for these, the bi-infinite linking sequence $(S_{k}^{x,y})$ satisfies:
$S_{k}^{x,y}=\begin{cases}x,&\text{if $k>0$}\\ y,&\text{if $k\leq 0$}\end{cases}$
Thus, for this bi-infinite sequence, we have
$\sum_{k\in\mathbb{Z}}[2^{-k\theta}\cdot J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y})]^{q}=[J_{M}(1;x,y)]^{q}.$
From this, it follows that
$p_{\theta,q}(x,y)\leq\Gamma_{\theta,q}((J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y}))\leq J_{M}(1;x,y).$ (13)
From inequality 13, we have that $p_{\theta,q}(x,y)<+\infty$ for all $x,y\in X_{0}\cap X_{1}$ and from inequality 12 we have that $p_{\theta,q}(x,y)=0\Longleftrightarrow x=y$. ∎
Lemma 8.
Let $\theta\in(0,1)$ and $q\in[1,\infty]$. Given $x,y\in X_{0}\cap X_{1}$, let us define the functional
$P_{\theta,q}(x,y):=\frac{p_{\theta,q}(x,y)+p_{\theta,q}(y,x)}{2}.$ (14)
The functional $P_{\theta,q}:(X_{0}\cap X_{1})\times(X_{0}\cap X_{1})\to[0,\infty)$ satisfies
  1. 1.
    $P_{\theta,q}(x,y)=0\Longleftrightarrow x=y$.
  2. 2.
    $P_{\theta,q}(x,y)=P_{\theta,q}(y,x)$, for all $x,y\in X_{0}\cap X_{1}$.
  3. 3.
    $\min\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}\leq P_{\theta,q}(x,y)\leq\max\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}$, for all $x,y\in X_{0}\cap X_{1}$.
Proof.
The proof is immediate. ∎
Definition 3.3.
Given $\theta\in(0,1)$ and $q\in[1,\infty]$ let us define the functional $\delta_{\theta,q}$ on $X_{0}\cap X_{1}$, given by
$\delta_{\theta,q}(x,y):=\inf\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1}),$ (15)
where the infimum is taken over all finite sequences of points $x_{0}=x,x_{1},x_{2},x_{3},\ \ldots,x_{n}=y\in X_{0}\cap X_{1}$.
Theorem 3.2.
The functional $\delta_{\theta,q}$ as above is a metric on $X_{0}\cap X_{1}$.
Proof.
It is easy to verify that $\delta_{\theta,q}(x,x)=0$. Furthermore, for all $x,y\in X_{0}\cap X_{1}$, taking the sequence $x_{0}=x$ and $x_{1}=y$, we have
$0\leq\delta_{\theta,q}(x,y)\leq P_{\theta,q}(x,y)<+\infty.$
Let now $x,y\in X_{0}\cap X_{1}$ such that $\delta_{\theta,q}(x,y)=0$. From inequality 12, we have
$K_{M}(1;x,y)=K_{M}(1;y,x)\leq M_{\theta,q}\cdot p_{\theta,q}(x,y),$
for all $x,y\in X_{0}\cap X_{1}$. Thus, it is also true that $K_{M}(1;x,y)\leq M_{\theta,q}\cdot p_{\theta,q}(y,x)$. Therefore,
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot\min\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}\leq M_{\theta,q}\cdot P_{\theta,q}(x,y),$ (16)
for any $x,y\in X_{0}\cap X_{1}$. For any linking sequence $x_{0}=x,x_{1},x_{2},\ldots,x_{n}=y$ we have
$K_{M}(1;x,y)$ $\leq\sum_{j=0}^{n-1}K_{M}(1;x_{j},x_{j+1})$
$\leq M_{\theta,q}\cdot\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1}).$
Consequently, we have
$K_{M}(1;x,y)\leq M_{\theta,q}\cdot\delta_{\theta,q}(x,y),\quad x,y\in X_{0}\cap X_{1}.$ (17)
This inequality guarantees the positivity of $\delta_{\theta,q}$. The symmetry of $\delta_{\theta,q}$ comes from the symmetry of $P_{\theta,q}$ and the fact that every linking sequence from $x$ to $y$ in $X_{0}\cap X_{1}$ is also a linking sequence from $y$ to $x$ in $X_{0}\cap X_{1}$. For the triangle inequality, let $x,y,z\in X_{0}\cap X_{1}$ and $\varepsilon>0$. Then, there exist linking sequences $x_{0}=x,x_{1},x_{2},\ldots,x_{n}=z$ from $x$ to $z$ and $y_{0}=z,y_{1},y_{2},\ldots,y_{m}=y$ from $z$ to $y$, such that
$\delta_{\theta,q}(x,z)\leq\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1})<\delta_{\theta,q}(x,z)+\varepsilon$
and
$\delta_{\theta,q}(z,y)\leq\sum_{j=0}^{m-1}P_{\theta,q}(y_{j},j_{j+1})<\delta_{\theta,q}(z,y)+\varepsilon.$
Note that the sequence
$z_{0}=x,z_{1},\ldots,z_{n}=z,z_{n+1}=z,z_{n+2},\ldots,z_{n+m}=y$
is a linking sequence from $x$ to $y$ in $X_{0}\cap X_{1}$ and satisfies
$\delta_{\theta,q}(x,y)$ $\leq\sum_{j=0}^{m+n-1}P_{\theta,q}(z_{j},z_{j+1})$
$=\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1})+\sum_{j=0}^{m-1}P_{\theta,q}(y_{j},y_{j+1})$
$<\delta_{\theta,q}(x,z)+\delta_{\theta,q}(z,y)+2\varepsilon.$
From the arbitrariness of $\varepsilon>0$ it follows
$\delta_{\theta,q}(x,y)\leq\delta_{\theta,q}(x,z)+\delta_{\theta,q}(z,y).$
Therefore, $\delta_{\theta,q}$ is indeed a metric on $X_{0}\cap X_{1}$. ∎
Remark 2.
Inequality 17 ensures that the metric space $(X_{0}\cap X_{1},\delta_{\theta,q})$ is continuously included in the metric space $(X_{0}\cup X_{1},K_{M}(1))$, thus satisfying the conditions for relative completion.
Definition 3.4 (The $R$-interpolated metric space).
The $R$-interpolated metric space is denoted by $R_{\theta,q}^{\overrightarrow{X}}$ and is the relative completion of the metric space
$(X_{0}\cap X_{1},\delta_{\theta,q})$
in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. The metric of this space is denoted by $\Delta_{\theta,q}$ and equals $\Delta_{\theta,q}(x,y)=\delta_{\theta,q}(x,y)$, for all $x,y\in X_{0}\cap X_{1}$.
Remark 3.
An important fact is that the interpolated spaces above were defined as the completion of the intersection $X_{0}\cap X_{1}$ with a specific metric in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. This ensures that the elements of the interpolated spaces are contained in $X_{0}\cup X_{1}$. However, note that such metrics on the intersection also satisfy the conditions for completion in $(X,d_{X})$ which can be substantially “larger” than $X_{0}\cup X_{1}$, providing many more new elements to the completions. That is why the name “relative” was chosen for this type of completion. We do not study here the advantages of completing the spaces beyond $X_{0}\cup X_{1}$.
4 Properties of the $R$-Interpolated Metric Space
Theorem 4.1.
The inclusions $X_{0}\cap X_{1}\subseteq R_{\theta,q}^{\overrightarrow{X}}\subseteq X_{0}\cup X_{1}$ are valid and furthermore
  1. 1.
    $\Delta_{\theta,q}(x,y)=\delta_{\theta,q}(x,y)\leq J_{M}(1;x,y),\quad x,y\in X_{0}\cap X_{1}$.
  2. 2.
    $K_{M}(1;x,y)\leq M_{\theta,q}\cdot\Delta_{\theta,q}(x,y),\quad x,y\in R_{\theta,q}^{\overrightarrow{X}}$.
Proof.
The inclusions are immediate. We will show only the inequalities. Let $x,y\in X_{0}\cap X_{1}$. From inequality 13, we have
$p_{\theta,q}(x,y)\leq J_{M}(1;x,y).$
Since $J_{M}(1;x,y)=J_{M}(1;y,x)$, it holds that
$\max\{p_{\theta,q}(x,y);p_{\theta,q}(y,x)\}\leq J_{M}(1;x,y).$
Thus, $P_{\theta,q}(x,y)\leq J_{M}(1;x,y)$. For the trivial linking sequence $x_{0}=x$ and $x_{1}=y$, we have
$\delta_{\theta,q}(x,y)\leq P_{\theta,q}(x,y)\leq J_{M}(1;x,y).$ (18)
The second inequality is obtained directly from inequality 17. Therefore, the inclusions are continuous, as we wanted to demonstrate. ∎
Definition 4.1.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces and $T:X\to Y$ an operator. We say that $T$ is a closed operator if
$d_{X}(x_{n},x)\to 0\quad\text{and}\quad d_{Y}(T(x_{n}),y)\to 0]\implies y=T(x).$
In other words, the set
$\text{Gr}(T):=\{(x,T(x))\in X\times Y\}$
is a closed set in $X\times Y$ in the product topology.
Definition 4.2.
Let $(X,d_{X})$, $(y,d_{Y})$ and $T:X\to Y$ an operator. We say that the operator $T$ is a Lipschitz operator if there exists a positive constant $C$ such that
$d_{Y}(T(x),T(y))\leq C\cdot d_{X}(x,y),\quad\forall x,y\in X.$
When $T$ is a Lipschitz operator, we call the number
$\inf_{C>0}\{d_{Y}(T(x),T(y))\leq C\cdot d_{X}(x,y);\quad\forall x,y\in X\}$
the Lipschitz constant of the operator $T$.
Lemma 9 (Reindexing and interpolation factor).
Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces and $T:X\to Y$ an operator. Let $(S_{k})_{k\in\mathbb{Z}}$ be a linking sequence from $x$ to $y$ in $X$ (which for simplicity of notation we will omit the superscript $x,y$) and let
$a_{k}:=2^{-k\theta}J_{M}(2^{k}\lambda;\,T(S_{k}),T(S_{k+1}):Y),\qquad b_{k}:=2^{-k\theta}J_{M}(2^{k};\,S_{k},S_{k+1}:X),$
where $\lambda>0$ and $\theta\in(0,1)$. Suppose that for all $k\in\mathbb{Z}$ it holds that
$J_{M}(2^{k};S_{k},S_{k+1}:X)\;\geq\;\omega_{0}^{-1}\,J_{Y}(2^{k}\lambda;\,T(S_{k}),T(S_{k+1}):Y),\quad\omega_{0}>0.$ (H)
For $\omega_{1}>0$ define $\lambda=\dfrac{\omega_{0}}{\omega_{1}}$. Then:
  1. (i)
    If $\lambda=2^{r}$ for some $r\in\mathbb{Z}$, then, for all $q\in[1,\ infty]$,
    $\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\;\big\|(b_{k})_{k}\big\|_{\ell^{q}}.$
  2. (ii)
    If $\lambda>0$ is arbitrary, then, for all $q\in[1,\infty]$,
    $\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq C(\theta)\,\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\;\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
    where $C(\theta)\leq 2^{\theta}$ is a constant depending only on $\theta$.
Proof.
We start by observing that hypothesis (H) is equivalent, in terms of the terms $a_{k},b_{k}$, to
$b_{k}\;\geq\;\omega_{0}^{-1}a_{k},\qquad\text{that is}\qquad a_{k}\leq\omega_{0}\,b_{k},\qquad\forall k\in\mathbb{Z}.$
If $\lambda=2^{r}$, $r\in\mathbb{Z}$: Writing $\lambda=2^{r}$ and making the index change $m=k+r$ we obtain, for each $k$,
$a_{k}=2^{-k\theta}J_{M}(2^{k}\lambda;\,T(S_{k}),T(S_{k+1}):Y)=2^{-k\theta}J_{M}(2^{k+r};\,T(S_{k}),T(S_{k+1}):Y).$
Rewriting in terms of the index $m$ (with $m=k+r$) we have the identity
$a_{k}$ $=2^{r\theta}\,2^{-(k+r)\theta}J_{M}(2^{k+r};\,T(S_{k}),T(S_{k+1}):Y)$
$=\lambda^{\theta}\,2^{-m\theta}J_{M}(2^{m};\,T(S_{m-r}),T(S_{m-r+1}):Y).$
By hypothesis (H) applied to the index mrm-r we have
$2^{-m\theta}J_{M}(2^{m};\,T(S_{m-r}),T(S_{m-r+1}):Y)\leq$
$\omega_{0}\,2^{-(m-r)\theta}J_{M}(2^{m-r};\,S_{m-r},S_{m-r+1}:X).$
Combining the two expressions and using reindexing (the transformation $k\mapsto m=k+r$ is an index shift), we obtain
$2^{-m\theta}J_{M}(2^{m};\,T(S_{m}),T(S_{m+1}):Y)\leq$
$\omega_{0}\,\lambda^{-\theta}\,2^{-(m-r)\theta}J_{M}(2^{m-r};\,S_{m-r},S_{m-r+1}:X).$
Taking the $\ell^{q}$ norm of both sides and remembering that the index shift is isometric in $\ell^{q}$ (that is, $\|(u_{m-r})_{m}\|_{\ell^{q}}=\|(u_{k})_{k}\|_{\ell^{q}}$), it results in
$\big\|(2^{-m\theta}J_{M}(2^{m};T(S_{m}),T(S_{m+1});Y))_{m}\big\|_{\ell^{q}}\leq$
$\omega_{0}\,\lambda^{-\theta}\,\big\|(2^{-k\theta}J_{X}(2^{k};S_{k},S_{k+1}:X))_{k}\big\|_{\ell^{q}}.$
But the left-hand side is exactly $\|(a_{k})_{k}\|_{\ell^{q}}$ (after reindexing), and the right-hand side is $\|(b_{k})_{k}\|_{\ell^{q}}$. Substituting $\lambda=\dfrac{\omega_{0}}{\omega_{1}}$ we obtain
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}\Big(\frac{\omega_{0}}{\omega_{1}}\Big)^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}}=\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
which proves (i).
Arbitrary Case $\lambda>0$: Choose $r=\lfloor\log_{2}\lambda\rfloor\in\mathbb{Z}$ such that $2^{r}\leq\lambda<2^{r+1}$. For all $k$ we have the two elementary estimates
$J_{M}(2^{k+r};\cdot,\cdot:Y)\;\leq\;J_{M}(2^{k}\lambda;\cdot,\cdot:Y)\;\leq\;J_{Y}(2^{k+r+1};\cdot,\cdot:Y),$
because the function $s\mapsto J_{M}(2^{k}s;\cdot,\cdot:Y)$ is non-decreasing in $s>0$ (remembering that here only the multiplicative parameter within the second argument of $J_{M}$ varies). Multiplying by $2^{-k\theta}$ and reindexing as before, we obtain
$2^{r\theta}\,2^{-(k+r)\theta}J_{M}(2^{k+r};\cdot,\cdot:Y)$ $\leq 2^{-k\theta}J_{M}(2^{k}\lambda;\cdot,\cdot:Y)$
$\leq 2^{(r+1)\theta}\,2^{-(k+r+1)\theta}J_{M}(2^{k+r+1};\cdot,\cdot:Y).$
Applying inequality (H) (in the form used in (i)) to the appropriate indices and taking the $\ell^{q}$ norm, we arrive at inequalities of the type
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}\,2^{\theta}\,\lambda^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}}\quad\text{or}\quad\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq\omega_{0}\,\lambda^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
depending on whether we use the majorant or the minorant above. In both cases we can write compactly
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq C(\theta)\,\omega_{0}\,\lambda^{-\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},\qquad C(\theta)\leq 2^{\theta}.$
Substituting $\lambda=\dfrac{\omega_{0}}{\omega_{1}}$ results in
$\big\|(a_{k})_{k}\big\|_{\ell^{q}}\leq C(\theta)\,\omega_{0}^{\,1-\theta}\omega_{1}^{\,\theta}\big\|(b_{k})_{k}\big\|_{\ell^{q}},$
as we wanted to demonstrate. ∎
Theorem 4.2.
Let $(X_{0},d_{X_{0}})$ and $(X_{1},d_{X_{1}})$ be a pair of compatible metric spaces with respect to the metric space $(X,d_{X})$ and $(Y_{0},d_{Y_{0}})$ and $(Y_{1},d_{Y_{1}})$ a pair of compatible metric spaces with respect to the metric space $(Y,d_{Y})$. If $(Y,d_{Y})$ is complete and $T:X\to Y$ is a closed operator that satisfies:
  1. 1.
    $T_{0}:=T|_{X_{0}}:X_{0}\to Y_{0}$ is Lipschitzian with Lipschitz constant $\omega_{0}>0$
  2. 2.
    $T_{1}:=T|_{X_{1}}:X_{1}\to Y_{1}$ is Lipschitzian with Lipschitz constant $\omega_{1}>0$
Then, $T_{\theta,q}:=T|_{R_{\theta,q}^{\overrightarrow{X}}}:R_{\theta,q}^{\overrightarrow{X}}\to R_{\theta,q}^{\overrightarrow{Y}}$ is Lipschitzian with Lipschitz constant not exceeding $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}>0$.
Proof.
Let $x,y\in X_{0}\cap X_{1}$ and $(x_{n})$ and $(y_{n})$ be two approximating sequences of $x$ and $y$ in $X_{0}\cap X_{1}$. Let us define the bi-infinite linking sequence $(S_{k}^{x,y})$ that links $x$ to $y$ in $X_{0}\cap X_{1}$ as usual through these. Since $(x_{n})$ is Cauchy in $(X_{0}\cap X_{1},J_{M}(1))$, it follows that $(x_{n})$ is Cauchy in $(X_{0},d_{X_{0}})$ and in $(X_{1},d_{X_{1}})$, because
$J_{M}(1;x_{m},x_{n}:X)$ $:=\max\{d_{X_{0}}(x_{m},x_{n}),d_{X_{1}}(x_{m},x_{n})\}$
$\geq d_{X_{i}}(x_{m},x_{n}),\quad i=0,1.$
Since $(x_{n})$ is Cauchy in $(X_{i},d_{X_{i}})$, $i=0,1$, due to the fact that
$d_{Y_{i}}(T_{i}(x_{m}),T_{i}(x_{n}))\leq\omega_{i}\cdot d_{X_{i}}(x_{m},x_{n}),\quad i=0,1,$
we have that $(T(x_{n}))$ is a Cauchy sequence both in $(Y_{0},d_{Y_{0}})$ and in $(Y_{1},d_{Y_{1}})$. Since by hypothesis there exists $C_{i}>0$ such that $d_{Y}(T(x_{m}),T(x_{n}))\leq C_{i}\cdot d_{Y_{i}}(T(x_{m}),T(x_{n}))$, for $i=0,1$, we have that $(T(x_{n}))$ is a Cauchy sequence in $(Y,d_{Y})$, which is complete, by hypothesis. It follows that there exists $a\in Y$ such that $d_{Y}(T(x_{n}),a)\to 0$. By the closedness property of the operator $T$, we have $a=T(x)$. The same reasoning can be applied to the sequence $(y_{n})$ which will provide $b\in Y$ such that $T(y)=b$. Note that if $(T(S_{k}^{x,y}))$ is a bi-infinite linking sequence from $T(y)$ to $T(x)$ in $Y_{0}\cap Y_{1}$. Thus, since
$J_{M}(2^{k};S_{k}^{x,y},S_{k+1}^{x,y}:X)=$
$\max\{d_{X_{0}}(S_{k}^{x,y},S_{k+1}^{x,y});2^{k}d_{X_{1}}(S_{k}^{x,y},S_{k+1}^{x,y})\}\geq$
$\max\{\omega_{0}^{-1}d_{Y_{0}}(T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)}),2^{k}\omega_{1}^{-1}d_{Y_{1}}(T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)})\}=$
$\omega_{0}^{-1}\cdot J_{M}(2^{k}\omega_{0}/\omega_{1};T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)}):Y)\geq$
$2^{-\theta}\cdot\omega_{0}^{\theta-1}\cdot\omega_{1}^{-\theta}\cdot J_{M}(2^{k};T(S_{k}^{T(x),T(y)}),T(S_{k+1}^{T(x),T(y)}):Y),$
where in the last inequality we used the previous Lemma 9. From this inequality, we have that
$p_{\theta,q;Y}(T(x),T(y))\leq 2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}\cdot p_{\theta,q;X}(x,y).$
From this inequality, it is immediate that
$\Delta_{\theta,q;Y}(T(x),T(y))\leq 2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in R_{\theta,q}^{\overrightarrow{X}}$ (19)
And this concludes the proof. ∎
Remark 4.
If we use the result that
$J_{M}(t;\cdot,\cdot)\leq\max\{1,t/s\}\cdot J_{M}(s;\cdot,\cdot)$
we can eliminate the dependence on $\theta$ from the upper bound of the Lipschitz constant of the operator $T$ in the interpolated space. In this situation, setting $t=2^{k}$ and $s=2^{k}\omega_{0}/\omega_{1}$, we have
$\max\{\omega_{0},\omega_{1}\}\cdot J_{M}(2^{k};S_{k},S_{k+1}:X)\geq J_{M}(2^{k};T(S_{k}),T(S_{k+1}):Y),$
resulting in
$\Delta_{\theta,q;Y}(T(x),T(y))\leq\max\{\omega_{0},\omega_{1}\}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in R_{\theta,q}^{\overrightarrow{X}}.$ (20)
We choose to display the possible dependence on $\theta$ because it is a constant more characteristic of the language in Interpolation Theory and expresses some type of convexity.
Corollary 1.
In Theorem 4.2 consider $X=Y$ and $d_{X}=d_{Y}$ a metric space complete. If $\omega_{i}<1$ for $i=0,1$, and $(X_{0},d_{X_{0}})$ and $(X_{1},d_{X_{1}})$ are both complete then there exists a unique fixed point of $T$ in $X_{0}\cup X_{1}$, even if $T$ is not necessarily Lipschitzian nor continuous in $(X_{0}\cup X_{1},K_{M}(1))$.
Proof.
Since $\max\{\omega_{0},\omega_{1}\}<1$, it follows from inequality 20 that $T_{\theta,q}$ (the operator $T$ restricted to the interpolated space) has a unique fixed point $x^{\prime}$, by the Banach Fixed Point Theorem. Such a fixed point necessarily belongs to $X_{0}\cup X_{1}$. Thus, $x^{\prime}\in X_{0}$ or $x^{\prime}\in X_{1}$. Since by hypothesis the operator $T$ is a strong contraction in $X_{0}$ and in $X_{1}$, unique fixed points $x_{0}\in X_{0}$ and $x_{1}\in X_{1}$ exist. Therefore, we must have $x^{\prime}=x_{0}=x_{1}$ and no other distinct point can be fixed by $T$ in $X_{0}\cup X_{1}$. ∎
Remark 5.
Throughout the proof we used the same letter $T$ to indicate different operators, since these operators were only restrictions of the same $T$ to spaces that were clear in the context through other symbols.
Remark 6.
In Theorem 4.2 if we impose the condition of continuity on $T:X\to Y$ we can waive the completeness of $(Y,d_{Y})$.
Definition 4.3.
Let $T:X\to Y$ be an operator between metric spaces. We will say that $T$ is a compact operator if for every sequence $(x_{n})$ of points in $X$ that is bounded, there exists a subsequence $(x_{n_{k}})$ of $(x_{n})$ such that $(T(x_{n_{k}}))$ is convergent.
Proposition 4.1.
If $X_{0}=X_{1}=X$ and $d_{0}=d_{1}=d_{X}$. Given $\theta\in(0,1)$ and $q\in[1,\infty]$, we have
$R_{\theta,q}^{\overrightarrow{X}}=X$
and the metric $\Delta_{\theta,q}$ is equivalent to the metric $d_{X}$.
Proof.
For $t>0$ and $u,v\in X$,
$J_{M}(t;u,v)=\max\{d_{X}(u,v),\,t\cdot d_{X}(u,v)\}=\begin{cases}d_{X}(u,v),&0<t\leq 1,\\ t\cdot d_{X}(u,v),&t\geq 1.\end{cases}$
In particular, for $t=1$ we have $J_{M}(1;u,v)=d_{X}(u,v)$, and the metric $K_{M}(1;\cdot,\cdot)$ defined by infimum over linking sequences coincides with the original metric $d_{X}$ (since the cost per step is always $h_{1}(\cdot,\cdot)=d_{X}(\cdot,\cdot)$). By inequalities (12) and (13)):
$K_{M}(1;u,v)\leq M_{\theta,q}\cdot p_{\theta,q}(u,v),$ (21)
and for constant sequences,
$p_{\theta,q}(u,v)\leq J_{M}(1;u,v).$ (22)
In the trivial case, $K_{M}(1;u,v)=d_{X}(u,v)$ and $J_{M}(1;u,v)=d_{X}(u,v)$. Substituting these equalities into (21) and (22) we obtain, for all $u,v\in X$,
$d_{X}(u,v)\leq M_{\theta,q}\cdot p_{\theta,q}(u,v)\qquad\text{and}\qquad p_{\theta,q}(u,v)\leq d_{X}(u,v).$
Rewriting,
$\frac{1}{M_{\theta,q}}\cdot d_{X}(u,v)\;\leq\;p_{\theta,q}(u,v)\;\leq\;d_{X}(u,v).$ (23)
The inequalities (23) show that $p_{\theta,q}$ is equivalent to the metric $d_{X}$ (with explicit constants $1/M_{\theta,q}$ and $1$). We have
$P_{\theta,q}(u,v)=\tfrac{1}{2}\bigl(p_{\theta,q}(u,v)+p_{\theta,q}(v,u)\bigr),$
and
$\delta_{\theta,q}(x,y)=\inf\sum_{j=0}^{n-1}P_{\theta,q}(x_{j},x_{j+1}),$
where the infimum runs over all linking sequences $x_{0}=x,\dots,x_{n}=y$ in $X_{0}\cap X_{1}=X$. Since $p_{\theta,q}$ satisfies (23), the same inequalities hold for $P_{\theta,q}$. More precisely, using (23) and the definition of $P_{\theta,q}$, we obtain
$\frac{1}{M_{\theta,q}}\cdot d_{X}(u,v)\leq P_{\theta,q}(u,v)\leq d_{X}(u,v),\qquad\forall u,v\in X.$
Now, by the definition of $\delta_{\theta,q}$, we have
$\frac{1}{M_{\theta,q}}\cdot d_{X}(x,y)\leq\delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ inX.$
Therefore $\delta_{\theta,q}$ is equivalent to $d_{X}$. Since $R_{\theta,q}^{\overrightarrow{X}}$ was defined as the relative completion of $(X_{0}\cap X_{1},\delta_{\theta,q})$ in $(X_{0}\cup X_{1},K_{M}(1))$ and here $X_{0}\cap X_{1}=X_{0}\cup X_{1}=X$, it holds that
$R_{\theta,q}^{\overrightarrow{X}}=X.$
Since $\Delta_{\theta,q}=\delta_{\theta,q}$ on $X_{0}\cap X_{1}=X$, we have
$\frac{1}{M_{\theta,q}}\cdot d_{X}(x,y)\leq\Delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ \in X$
as we wanted to demonstrate. ∎
Corollary 2.
Let $(X_{0},X_{1})_{X}$ be a pair of metric spaces, $\theta\in(0,1)$, $q\in[1,\infty]$ and $(F,d_{F})$ a complete metric space. If $T$ is an operator that satisfies
  • (i)
    $T:X\rightarrow F$ is closed;
  • (ii)
    $T:X_{0}\rightarrow F$ is a compact Lipschitzian operator;
  • (iii)
    $T:X_{1}\rightarrow F$ is Lipschitzian;
  • (iv)
    Every sequence bounded in $X_{0}\cap X_{1}$ with respect to the interpolated metric is also bounded with respect to the metric $d_{X_{0}}$,
then $T_{\theta,q}:=T|_{R_{\theta,q}^{\overrightarrow{X}}}:R_{\theta,q}^{\overrightarrow{X}}\to F$ is Lipschitzian and compact.
Proof.
The result is immediate from the definitions, proposition 4.1 and theorem 4.2. ∎
Corollary 3.
Let $(X_{0},X_{1})_{X}$ be a pair of compatible metric spaces, $\theta\in(0,1)$, $q\in[1,\infty)$ and $T:X\to X$ a closed operator that satisfies
  • (a)
    $T_{0}:X_{0}\rightarrow X_{0}$ is a lipschitzian operator with constant $\omega_{0}$;
  • (b)
    $T_{1}:X_{1}\rightarrow X_{1}$ is a lipschitzian operator with constant $\omega_{1}$;
  • (c)
    $X_{1}$ is closed in $(X_{0}\cup X_{1},K_{M}(1))$ and $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}<1$.
Then $T_{1}$ has a fixed point.
Proof.
Due to Theorem 4.2, the operator $T$ restricted to the interpolated space $R_{\theta,q}^{\overrightarrow{X}}$ is lipschitzian with constant not exceeding $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}$. Thus, $T_{\theta,q}$ is a strong contraction. Since $(X,d_{X})$ is complete, it follows that the relative completion is also complete. Thus, by Banach’s Fixed-Point Theorem, the operator $T_{\theta,q}$ has a unique fixed point $w$ in $R_{\theta,q}^{\overrightarrow{X}}$. By the density of $X_{0}\cap X_{1}$ in the interpolated space and the continuous inclusion of the interpolated space in the metric space $(X_{0}\cup X_{1},K_{M}(1))$, there exists a sequence $(w_{n})$ of points in $X_{0}\cap X_{1}$ such that $K_{M}(1;w_{n},w)\to 0$. Since $X_{1}$ is closed in $(X_{0}\cup X_{1},K_{M}(1))$, it follows that $w\in X_{1}$. Thus, due to the closedness of $T$ in all $X$, it follows that $T_{1}(w)=w$.
5 The $R$-method in Banach Spaces
Let $(X_{0},\|\cdot\|_{0})$ and $(X_{1},\|\cdot\|_{1})$ be Banach spaces continuously embedded into a common Hausdorff topological vector space $Z$ over a field $\mathbb{F}$ ($\mathbb{R}$ or $\mathbb{C}$), with $X_{0}\cap X_{1}\neq\emptyset$. Let us consider the classic $K$-functional due to Lions-Peetre[see [10]] given by
$K(t;x):=\inf_{x_{0}+x_{1}=x}(\|x_{0}\|_{0}+t\cdot\|x_{1}\|_{1}),\quad t>0,\quad x\in X_{0}+X_{1}.$ (24)
It is known that for all $t>0$ the functional $K(t;\cdot)$ is a norm on $X=X_{0}+X_{1}$ and that $(X_{0}+X_{1},K(t;\cdot))$ is a Banach space.
Note that $K(1;x)\leq\|x\|_{0}$ for all $x\in X_{0}$ and $K(1;y)\leq\|y\|_{1}$ for all $y\in X_{1}$. It shows that $(X_{0},X_{1})_{X}$ is a compatible pair of metric spaces. In this environment, we have
Proposition 5.1.
$R_{\theta,q}^{\overrightarrow{X}}$ is a Banach space.
Proof.
Note that $0\in X_{0}\cap X_{1}\subseteq R_{\theta,q}^{\overrightarrow{X}}$. Now, we define $\|x\|_{\theta,q}:=\Delta_{\theta,q}(x,0)$ for all $x\in R_{\theta,q}^{\overrightarrow{X}}$. The set $X_{0}\cap X_{1}$ is endowed with a linear structure. By density of $X_{0}\cap X_{1}$ we can define a linear structure in $R_{\theta,q}^{\overrightarrow{X}}$ by putting for all $x,y\in R_{\theta,q}^{\overrightarrow{X}}$ and $\lambda\in\mathbb{F}$
$x+\lambda\cdot y:=\lim_{n\to\infty}x_{n}+\lambda\cdot y_{n},$ (25)
where $\|x_{n}-x\|_{\theta,q}\to 0$ and $\|y_{n}-y\|_{\theta,q}\to 0$ and $(x_{n})$ and $(y_{n})$ are sequences of points in $X_{0}\cap X_{1}$ given by the density. By the equality 25, it follows that $\|\cdot\|_{\theta,q}$ is a norm on $R_{\theta,q}^{\overrightarrow{X}}$ because $\delta_{\theta,q}=\Delta_{\theta,q}$ in $X_{0}\cap X_{1}$. Furthermore, since $(X_{0},\|\cdot\|_{0})$ and $(X_{1},\|\cdot\|_{1})$ are Banach spaces, it is also known that $X=X_{0}+X_{1}$ is a Banach space with the norm $K(1;\cdot)$. Thus, $R_{\theta,q}^{\overrightarrow{X}}$ is a complete metric space, because it is the relative completion of $(X_{0}\cap X_{1},\|\cdot\|_{\theta,q})$ in $(X_{0}+X_{1},K(1;\cdot))$ and therefore a Banach Space.
Remark 7.
This feature highlights a genuine advantage of the $R$-method. Classical interpolation functors on normed spaces are formulated within a linear framework and act exclusively on linear continuous operators. By contrast, the $R$-method is intrinsically metric and does not rely on linear structure. When restricted to normed spaces, it still induces an interpolation functor and preserves the Lipschitz property of operators with closed graphs, allowing for the interpolation of nonlinear mappings.
In particular, the method enlarges the classical category of admissible operators while retaining the functorial behavior expected in interpolation theory, thereby providing a natural extension of classical interpolation beyond the linear setting.